Diego is 1 year younger than Suzie. Trent is 19. Harold is 5 years older than Trent and 4 years younger than Suzie. How old is Suzie?

Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
;
Diego is 1 year younger than Suzie.
d = s - 1
:
Trent is 19. Harold is 5 years older than Trent
h = 19 + 5 = 24
:
and 4 years younger than Suzie.
h = s - 4
replace h with 24
24 = s - 4
24 + 4 = s
s = 28 yrs is Susie